The following maximum problem is in standard form. Introduce slack variables and set up the initial simplex.
Maximize P = 2x^1 + 4x^2 + x^3

Subject to the constraints:
x^1 + x^2 + x^3 <= 12
x^1 ≥ 0
x^2 ≥ 0
x^3 ≥ 0
x^1 + x^2 <= 6
2x^1 - x^2 + 3x^3 <= 6

Answer by Edwin McCravy(20055)

The following maximum problem is in standard form. Introduce slack variables and set up the initial simplex.
Maximize P = 2x1 + 4x2 + x3
Subject to the constraints:
x1 + x2 + x3 <= 12
x1 ≥ 0
x2 ≥ 0
x3 ≥ 0
x1 + x2 <= 6
2x1 - x2 + 3x3 <= 6

x1 + x2 + x3 + s1 = 12
x1 + x2 + s2 = 6
2x1 - x2 + 3x3 + s3 = 6
P = 2x1 + 4x2 + x3

Write that as this system:

 1x1 + 1x2 + 1x3 + 1s1 + 0s2 + 0s3 + 0P = 12
 1x1 + 1x2 + 0x3 + 0s1 + 1s2 + 0s3 + 0P =  6
 2x1 - 1x2 + 3x3 + 0s1 + 0s2 + 1s3 + 0P =  6
-2x1 - 4x2 - 1x3 + 0s1 + 0s2 + 1s3 + 1P =  0

Then make this as the first tableau:

 x1 x2 x3  s1 s2  s3  P
_____________________________
| 1  1  1 | 1  0  0 | 0 | 12 |
| 1  1  0 | 0  1  0 | 0 |  6 |
| 2 -1  3 | 0  0  1 | 0 |  6 |
|----------------------------|
|-2 -4 -1 | 0  0  1 | 1 |  0 |
ŻŻŻŻŻŻŻŻŻŻŻŻŻŻŻŻŻŻŻŻŻŻŻŻŻŻŻŻŻ

The first step will be to select the most negative
indicator which is -4. so the pivot column is the
2nd column. 

Divide each positive number above the -4 into the
corresponding element in the far right:
  
 ___         __ 
1)12 = 12,  1)6 = 6, 6 is smaller so pivot on the 1
in the 2nd row 2nd column. 

Edwin